OpenStudy (itsbribro):

Help A single tree produces about 2.6*10^2lb of oxygen each year. The amazon rain forest has about 3.9*10^11 trees. About how many ounces of oxygen are produced in the rain forest each day? 4.4*10^ ? oz

OpenStudy (triciaal):

1 tree = 2.6 *10^2 lb 3.9*10^11 trees = 2.6 *10^2lbs * 3.9*10^11 = x lbs 16 oz = 1 lb and you have x * 1lb so to convert that to ozs you would have 16 *x ozs do you follow?

OpenStudy (triciaal):

wait this is for the year the question asks to find for each day we just found how many ounces for the year and we know a year = 365 days can you find for 1 day?

OpenStudy (itsbribro):

im still really dont understand

OpenStudy (triciaal):

note that when doing conversions just need to keep the units in tact. example if you converted from year to day first you would get the same result

OpenStudy (triciaal):

read through the post again slowly from the top and let me know where you are stuck

OpenStudy (itsbribro):

@winsteria please help me!!!!!!!!

OpenStudy (winsteria):

ohgod well

OpenStudy (itsbribro):

i really dont undersatd at alll :(

OpenStudy (winsteria):

okay so what you need to do is.

OpenStudy (winsteria):

okay so one tree produces 2.6*10^2lb of oxygen a year so you need to take all of the trees that are in the forest which are 3.9*10^11 in all so then you need to divide the amount of all of the trees the forest by ohgod im stuck wait

OpenStudy (winsteria):

to be honest, i've never even done this math before but you need to divide the amount of trees in the forest by something (that something we don't know) and then multiply it by 2.6*10^2lb ???

OpenStudy (winsteria):

i'm sorry if i'm not much of any help but there are 365 days a year maybe divide that by the amount it produces a year??? idk

OpenStudy (itsbribro):

lol its ok do you know anyone who might know this?

OpenStudy (itsbribro):

and thank you so much

OpenStudy (winsteria):

i can tag some people for you and yw

OpenStudy (winsteria):

@ramiva @Roxeyrai73 can any of you help her with this problem, please she really needs help???

OpenStudy (itsbribro):

thank you so much

OpenStudy (anonymous):

i can't explain it other than how @triciaal did srry

OpenStudy (anonymous):

I would say that there are two ways to go about this, you can figure out how much the entire forest produces per year, and then divide by the days (365) to figure out how much it produces per day. Or you can divide the amount an individual tree generates per year by 365, to determine how much an individual tree generates per day, and then multiply that by the number of trees. you should end up with the same answer.

OpenStudy (itsbribro):

OMG I DONT KNOW

OpenStudy (anonymous):

so (2.6*10^2)(3.9*10^11) gives you the total amount of carbon that the amazon produces every year

OpenStudy (anonymous):

this gives you 1.014x10^14 lbs. per year

OpenStudy (anonymous):

to find lbs. per day we divide that number by the number of days per year which is 365

OpenStudy (anonymous):

this is approximately 2.78x10^11

OpenStudy (itsbribro):

ok so i would find the answer to 1.014x10^14 then divide the answer by 365?

OpenStudy (anonymous):

well 1.014x10^14 is just a value but its over 100 trillion so we write it in scientific notation to make it simple. so you just divide the value 1.014x10^14 by 365 to get the answer

OpenStudy (itsbribro):

okay well thank you for everything guys @winsteria @ramiva

OpenStudy (winsteria):

i wish i could have been of more help, i'm sorry D: but did you get the answer or???

OpenStudy (itsbribro):

yeah i figured it out

OpenStudy (itsbribro):

now im doing science :/ blah

OpenStudy (winsteria):

oh okay that's good and ikr science is just ugh ;-;

OpenStudy (itsbribro):

yeah it blows

OpenStudy (itsbribro):

im gonna message you so we can talk there and not on the thread

OpenStudy (winsteria):

okay